OpenStudy (awkwardpanda):

1. What is the slope of the line that passes through the given points? (6, -1) and (-3, -1) 2. Write the equation in slope-intercept form. What are the slope and y-intercept? -9x+10y=-9

OpenStudy (zale101):

m=y2-y1/x2-x1 for the first question

OpenStudy (zale101):

plug in the number and solve

OpenStudy (zale101):

numbers*

OpenStudy (anonymous):

So for #1: you could use\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] So you can choose which points to be x1,y1 and x2,y2, so if I chose (6,-1) as x1, y1 and (-3,-1) as x2,y2. It would be like\[m=\frac{ -1+1 }{ -3-6 }=\frac{ 0 }{ -9 }\]But as you can see here \[\frac{ 0 }{ 9 }\] is 0 since anything divided by 0 is 0. So you have a slope of 0. #2: So remember slope-intercept form is y = mx + b and the y-intercept is b and and slope is m. -9x+10y=-9 Solve for y, so add -9x to other side 10y = -9+9x Divide by 10\[y = \frac{ -9 }{ 10 }+\frac{ 9 }{ 10 }x\] so\[m=\frac{ 9 }{ 10 }\]and\[b=-\frac{ 9 }{ 10 }\] Hope this helps and Good Luck :)

OpenStudy (zale101):